# The interplay of the parametric instability and MRI in oscillatory shear flows

### v1

@@ 5.2 Interaction of MRI and PI
Having developed the base state MRI, we restart our simulation from the final snapshot output in section 5.1. As per our hydrodynamic sloshing run, we now switch on the sloshing forcing with a fiducial value of $\psi_{\rm f}=0.3$. We run the simulation for an additional $60T$ which tracks the growth of the PI on top of the MRI, and the eventual saturation into a quasi-steady sloshing state. The clear development of the PI is once again seen in Fig. 11.
@@
Figure 11: The emergence of the PI atop the MRI for our magnetized $\psi_{\rm f}=0.3$ run. Panel (a): the growth of $E_{\rm epi}$ is shown for the MHD run (black solid) compared with the hydro run (blue dotted) from section 4. Once again the laminar, resonant prediction is shown as the red dashed line. Panel (b): the rise in $\ln(E_{{\rm k},z})$ for the MHD (black solid) versus hydro (blue dotted) runs. The $y$-averaged radial sloshing flows $\{v_{x}\}_{y}$ at (c) $t=10T$ and (d) $t=60T$. The $y$-averaged vertical flows $\{v_{z}\}_{y}$ at (e) $t=10T$ and (f) $t=60T$.
@@
Panels (a) and (b) show the growth of $E_{\rm epi}$ and $\ln(E_{{\rm k},z})$, similarly to Fig. 5. Now, the new MHD simulation results are shown in black, whilst the previous hydrodynamic results are shown as the blue dotted line. The red dashed line shows the laminar resonant amplification of the sloshing motion, as before. Here we see that the initial epicylic growth is less pronounced. The MRI turbulence naturally resists the resonant growth and also seeds stronger fluctuations for the PI to latch onto. Together this reduces the initial growth of the epicylic motion and the PI takes over at earlier times – as indicated by the early rise phase of the vertical kinetic energy. Nonetheless, the maximum radial sloshing motions still reach an appreciable fraction of the sound speed and the ensuing PI has similar characteristics to the hydro run. Indeed, we see at late times that the epicylic energy drops down to a quasi-steady state, where the black and blue-dotted lines apparently overlap. This already hints that the final sloshing state is fairly similar in both the magnetized versus hydro setups. This is also supported by the velocity maps which are plotted for snapshots at $t=10T$ (c and e) and $t=60T$ (d and f). In panels (c) and (d) we plot $\{v_{x}\}_{y}$ which shows the clear presence of a sloshing flow which is broadly odd about the midplane, albeit with a distorted interface. This corrugation is once again driven by the emergence of a series of strong, counter-propagating vertical flows as visualized by $\{v_{z}\}_{y}$ in panels (e) and (f). The characteristic radial wavenumber in panel (d) is the same as that seen in Fig. 6, which is indicative of a similar, strong sloshing PI mechanism at play. However, we note that by the later snapshot (f), the radial wavenumber has decreased, indicating that the established elevator modes can still slowly evolve. Indeed, examining the intermediate times shows the merging of two channels, and the preference for this longer wavelength – perhaps aided by the MRI turbulence which diffuses and damps higher radial wavenumbers. Combining a model for the damping, facilitated by the MRI, with the PI Floquet growth rates for sufficiently subsonic sloshing amplitudes does indeed suggest that there is a preference for a longer wavelength to emerge at later times. This model will be discussed in more detail in section 5.4.
@@
Figure 12: Horizontal stresses in the equilibrium sloshing state, measured between $6040<t/T<$ 60. $\alpha_{\rm M}$ (red line) oscillates about the averaged value of $0.0099$ (dashed red). $\alpha_{\rm R}$ (blue line) oscillates about the averaged value of $0.0019$ (dashed blue).
@@
The sloshing state also modifies the nature of the horizontal stresses, compared with the standard MRI shown in Fig. 5.1. Note that when computing $\alpha_{\rm R}=\langle\rho u_{x}^{\prime}u_{y}^{\prime}\rangle_{xyz}/(q\langle\rho c^{2}\rangle_{xyz})$, the velocity fluctuations are now measured relative to the average sloshing flow. The results are shown in Fig. 12 over the time interval $6040<t/T<$ 60, after the sloshing flow has equilibrated. The Maxwell stresses $\alpha_{\rm M}$ (red) once again dominate over the Reynolds stresses $\alpha_{\rm R}$ (blue), with time-averaged values of $0.0099$ and $0.0019$ respectively. This corresponds to an amplification of $\sim 1.4$ compared with the base state. It should also be noted that the horizontal stresses exhibit a clear oscillation over the dynamical timescale, indicating that the horizontal fluctuations are also correlated with the forcing frequency.
@@ 5.3 MHD vertical stresses
We will now quantify the MHD equilibrium sloshing state in detail, akin to the analysis performed in section 4.4. In particular we will examine the balance between the sloshing flows and the vertical stresses, which now comprise both Reynolds and Maxwell terms. The Maxwell stresses are denoted
@@
and arise from the horizontal averaging of equation (2). In this way, we extend our horizontally-averaged equations for the sloshing motion by appending $-\partial\mathcal{M}_{xz}/\partial z$ and $-\partial\mathcal{M}_{yz}/\partial z$ to equations (27)-(28) respectively.
Figure 13: The rate-of-strain and stress relationship for our fiducial MHD run (akin to Fig. 9). Columns (1) and (2) respectively show the $x$ and $y$ components of the phase-folded, vertical profiles for the (a) sloshing motions, (b) Reynolds stresses and (c) Maxwell stresses, at different orbital phases indicated by the line color. Column (3) then extracts the corresponding amplitudes of these sinusoidal profiles as a function of the temporal phase.
In Fig. 13 we compare the amplitude and phase relationship between the equilibrium sloshing motion, $\mathcal{R}$ and $\mathcal{M}$. As per the analysis presented in Fig. 9, we phase fold our data modulo the dynamical period, over the interval $6040<t/T<$ 60, to boost our sloshing signal. In panels (a1) and (a2) we are plotting the horizontally-averaged sloshing motion $v_{x,{\rm w}}$ and $v_{y,{\rm w}}$, as a function of $z$, across a number of orbital phases. Similarly, we extract the phase-folded z profiles for $\mathcal{R}_{xz}$, $\mathcal{R}_{yz}$, $\mathcal{M}_{xz}$ and $\mathcal{M}_{yz}$ in panels (b1), (b2), (c1) and (c2). In keeping with the hydrodynamic results, we cleanly extract oscillating sinusoidal profiles for the sloshing motion, and $\cos(k_{\rm b}z)$ structures for the stresses. At each phase, we find the best fit to these profiles and thereby extract $A(t)$, $B(t)$ (panel (a3)), $R_{ij}$ (panel (b3)) and $M_{ij}$ (panel (c3)), where $\mathcal{M}_{ij}=M_{ij}(t)\cos{(k_{\rm b}z)}$.
By Fourier analyzing these curves, we isolate the amplitude and phase for the dominant orbital frequency component. We find that $|A|=0.221$ and $\arg(A)=-0.66$, whilst $|B|=0.079$ and $\arg(B)=2.06$. This shows that the sloshing amplitude is slightly larger in this case, compared with the previous hydrodynamic run. Performing the same analysis on the stresses we find $|R_{xz}|=0.0125$ and $\arg(R_{xz})=2.98$, whilst $|R_{yz}|=0.0134$ and $\arg(R_{yz})=-1.40$. Finally, $|M_{xz}|=0.0037$ and $\arg(M_{xz})=2.04$, whilst $|M_{yz}|=0.0031$ and $\arg(M_{yz})=-1.44$. Clearly the Reynolds stresses dominate the vertical transport, owing to the dominance of the PI in establishing strong elevator flows which resist the sloshing motions.
Once again we can find the viscoelastic linear relationship between the velocity shear and the stresses, such that $-\mathcal{R}_{ij}=\bar{\rho}\nu_{{\rm R},ij}\partial v_{i,\rm{w}}/\partial z$ and $-\mathcal{M}_{ij}=\bar{\rho}\nu_{{\rm M},ij}\partial v_{i,\rm{w}}/\partial z$. In addition to equation (30) we have an analogous expression involving the Maxwell viscoelastic coefficients
@@ 5.4 Variation with forcing amplitude
Figure 14: Variation in the sloshing diagnostics with $\psi_{\rm f}$. Row (a): Sloshing motion $A$ and $B$. Row (b): Reynolds stresses. Row (c): Maxwell stresses. Row (d): viscoelastic constants. Column (1) visualizes the full complex information in a polar plot whilst columns (2) and (3) plot the amplitude and phase information respectively, as a function of $\psi_{\rm f}$. The red and black line colors denote the $x$ vs $y$ variables whilst the different line-styles+markers denote the MHD, hydro and MHD-seeded hydro runs (see inset legend). The marker colors also denote the value of $\psi_{\rm f}$. Finally the starred markers correspond to the maximum radial sloshing amplitude attained during the transient phase after the forcing is switched on. For further explanation see the discussion in-text.
@@
Up until now we have focused on the fiducial forcing amplitude with $\psi_{\rm f}=0.3$. This value is sufficiently strong that the emerging sloshing equilibrium is well distinguished from the underlying MRI turbulence. However, in this section we will consider how the sloshing response varies as a function of $\psi_{\rm f}$. To this end we perform a suite of MHD simulations with the same setup as described in section 5.2 but with $\psi_{\rm f}\in\{0.01,0.02,0.03,0.05,0.08,0.1,0.2,0.3\}$. In all cases we first allow the sloshing to progress through the initial transient growth, before settling into a quasi-steady oscillation. We then monitor this between times $6040<t/T<$ 60 to extract the sloshing amplitudes, stresses and viscoelastic coefficients, as before. The results are shown in Fig. 14. Row (a) plots the extracted sloshing amplitudes, rows (b) and (c) show the Reynolds and Maxwell stresses respectively, and row (d) shows the viscoelastic constants. Column (1) visualizes the full complex amplitude and phase information in the form of polar plots. Meanwhile columns (2) and (3) break this up into the amplitude and phase variation as a function of $\psi_{\rm f}$. In all cases, the colored markers denote the different values for $\psi_{\rm f}$ whilst the solid, connective colored black and red lines with circular markers pertain to the ‘$x$’ versus ‘$y$’ quantities, as denoted by the inset labels. For comparison we also plot a sample of purely hydrodynamic runs in columns (2) and (3). Repeating the methodology described in section 4.1 for $\psi_{\rm f}\in\{0.01,0.03,0.1,0.3\}$, we similarly extract the sloshing diagnostics between $6040<t/T<$ 60 and plot the results as the dashed lines with triangular markers. Finally, we repeat these hydro runs but now start from the perturbed velocity and density initial condition inherited from the final snapshot of our base MRI run, with the magnetic field now switched off. These are shown as the dotted lines with square markers.
@@
Across all runs, panel (a2) clearly shows the expected increase in the sloshing response with the forcing amplitude. For lower values of $\psi_{\rm f}$, the ratio of $|A|/|B|\sim 2$ whilst the phase difference in panel (a3) is also roughly $\pi/2$, in keeping with a free epicylic motion. As $\psi_{\rm f}$ increases, this ratio $|A|/|B|$ rises and the phase difference trends towards $\pi$. Comparing the various physical setups, we clearly observe differences between the corresponding line styles. Notably the magnetized, MRI runs typically saturate with a higher sloshing amplitude compared with the hydro runs, particularly for intermediate values of $\psi_{\rm f}$. This is equivalently borne out by the magnetized runs showing lower viscoelastic amplitudes in panel (d2). This might seem surprising at first, since naively the action of the MRI might be thought to enhance the damping of the sloshing motions. However, the MRI turbulence also acts to damp the growth of the parametric instability and the associated elevator modes. Since the hydrodynamic, PI modes dominate the vertical stress, suppression of these via the MRI allows the sloshing to attain a larger amplitude. The fact that the hydro run, with the MRI inherited initial conditions, also deviates from the full MHD run, suggests that the final sloshing state is indeed sensitive to the different physics and not just the seeding of perturbations. However, we do notice that the MRI seeded hydro run does saturate at a larger sloshing amplitude for $\psi_{\rm f}=0.01$ compared with the standard, weakly seeded hydro setup. This might be explained by the earlier growth of the PI, which then taps the sloshing at an initially lower amplitude and hence sets up weaker elevator modes. These are then less effective at tempering the equilibrium sloshing.
@@
In all panels the vertical, dashed line and the bounded gray area, for $\psi_{\rm f}<\psi_{c}=0.03$, delimits the region for which the MRI is found to suppress the growth of the PI in our MHD runs. Namely, below this value the runs did not show the development of elevator modes and instead the turbulence resembles the MRI base state. Meanwhile, the elevator modes and PI are still found to emerge below this value in our purely hydrodynamic runs. This MHD suppression of the PI below $\psi_{c}$ also seems to manifest in the slopes of the sloshing and viscoelastic amplitudes. The sloshing amplitude in panel (a2) rises more quickly with $\psi_{\rm f}$ when the MRI suppresses the PI, before the ‘knee’ at $\psi_{c}$ where the response flattens. The viscoelastic coefficients also seem to rise more rapidly beyond this critical forcing as the PI overcomes the MRI.
@@
A simplified toy model might be invoked to motivate this critical forcing value. Specifically, the elevator modes can emerge when the PI growth rate, at the maximal value of sloshing during the initial transient phase, exceeds the damping induced by the turbulent transport of momentum across the boundary between elevator mode channels. This turbulent damping of the elevator modes can be estimated as $\sigma_{\rm d}\sim 4k_{x}v_{x,{\rm MRI}}/(2\pi)$. The factor of 4 corresponds to the momentum transport in/out of the two boundaries, either side of a vertical elevator channel. The $k_{x}=2\pi n_{x}/L_{x}$ encodes the radial width of the elevator mode, given the radial mode order $n_{x}$ fitting into the box of width $L_{x}$. Finally the $v_{x,{\rm MRI}}$ denotes a characteristic radial velocity associated with the MRI eddies. Here, we take this to be equal to the standard deviation of the velocity fluctuation distribution, $v_{x,{\rm MRI}}=0.085$, as found in Fig. 10. This competes with the PI growth rates $\sigma_{\rm PI}$, which can be found using the Floquet theory (using $N_{\rm max}=30$), as described in section 3.2. We solve for the maximally growing mode across a number of sloshing amplitudes $S$ and radial modes $n_{x}\leq 30$. Then for each value of $S$ we find the maximum net growth rate $\sigma_{\rm net}=\sigma_{\rm PI}-\sigma_{\rm d}$ and the corresponding optimal modal number $n_{x}$. The results are shown in Fig. 15 where the maximal growth rate is normalized by the value of $S$.
@@
Figure 15: A model for the competition between PI growth and MRI damping of elevator modes. The maximum net growth rate $\sigma_{\rm net}$, normalized by the sloshing amplitude $S$, is plotted as a function of $S$. The color of the markers denote the preferred radial order $n_{x}$ corresponding to this maximal growth rate.
@@
Furthermore, the color of the starred data points indicates the modal number $n_{x}$ with the largest growth rate. For $S\leq 0.1$ the MRI induced damping suppresses the PI growth for all modes. However, above this value our model predicts growth at $n_{x}=2$, before shorter wavelengths can overcome the damping at larger values of $S$. As one progresses to larger values of $S$ we see the preference for $n_{x}=5$, then $n_{x}=6$, in keeping with the elevator flows seen in our fiducial runs. As one progresses above $S\sim 1$ we find the optimal $n_{x}$ decreases again, in accordance with the leftward drifting peaks seen in Fig. 6.
@@
To draw comparison with the behavior seen in our MHD simulations, we return to Fig. 14 and plot the maximum radial, sloshing amplitude during the transient phase as starred markers in panel (a2). For $\psi_{\rm f}=0.03$ this yields a maximal sloshing of $|A|=0.12$, which should be susceptible to growth of the $n_{x}=2$ elevator mode according to Fig. 15. Upon examining the sloshing state for this run we do indeed see the emergence of this very mode – marking the transition to the PI dominated regime. Note that the maximal sloshing amplitude is always reduced substantially as the PI feeds off this shear energy and establishes stresses which equilibrate the sloshing at a lower amplitude.

### v2

@@ 5.2 Interaction of MRI and PI
Having developed the base state MRI, we restart our simulation from the final snapshot output in Section 5.1. As per our hydrodynamic sloshing run, we now switch on the sloshing forcing with a fiducial value of $\psi_{\rm f}=0.3$. We run the simulation for an additional $60T$, which tracks the growth of the PI on top of the MRI, and the eventual saturation into a quasi-steady sloshing state. The clear development of the PI is once again seen in Fig. 11.
@@
Figure 11: The emergence of the PI atop the MRI for our magnetized $\psi_{\rm f}=0.3$ run. Panel (a): The growth of $E_{\rm epi}$ is shown for the MHD run (black solid) compared with the hydrodynamic run (blue dotted) from Section 4. Once again, the laminar, resonant prediction is shown as the red dashed line. Panel (b): the rise in $\ln(E_{{\rm k},z})$ for the MHD (black solid) versus hydrodynamic (blue dotted) runs. The $y$-averaged radial sloshing flows $\{v_{x}\}_{y}$ at (c) $t=10T$ and (d) $t=60T$. The $y$-averaged vertical flows $\{v_{z}\}_{y}$ at (e) $t=10T$ and (f) $t=60T$.
@@
Panels (a) and (b) show the growth of $E_{\rm epi}$ and $\ln(E_{{\rm k},z})$, similarly to Fig. 5. Now, the new MHD simulation results are shown in black, whilst the previous hydrodynamic results are shown as the blue dotted line. The red dashed line shows the laminar resonant amplification of the sloshing motion, as before. Here we see that the initial epicylic growth is less pronounced. The MRI turbulence naturally resists the resonant growth and also seeds stronger fluctuations for the PI to latch onto. Together, this reduces the initial growth of the epicylic motion, and the PI takes over at earlier times – as indicated by the early rise phase of the vertical kinetic energy. Nonetheless, the maximum radial sloshing motions still reach an appreciable fraction of the sound speed, and the ensuing PI has similar characteristics to the hydrodynamic run. Indeed, we see at late times that the epicylic energy drops down to a quasi-steady state, where the black and blue dotted lines apparently overlap. This already hints that the final sloshing state is fairly similar in both the magnetized versus hydrodynamic setups. This is also supported by the velocity maps, which are plotted for snapshots at $t=10T$ ((c) and (e)) and $t=60T$ ((d) and (f)). In panels (c) and (d), we plot $\{v_{x}\}_{y}$, which shows the clear presence of a sloshing flow that is broadly odd about the midplane, albeit with a distorted interface. This corrugation is once again driven by the emergence of a series of strong, counter-propagating vertical flows as visualized by $\{v_{z}\}_{y}$ in panels (e) and (f). The characteristic radial wavenumber in panel (d) is the same as that seen in Fig. 6, which is indicative of a similar, strong sloshing PI mechanism at play. However, we note that by the later snapshot (f), the radial wavenumber has decreased, indicating that the established elevator modes can still slowly evolve. Indeed, examining the intermediate times shows the merging of two channels, and the preference for this longer wavelength – perhaps aided by the MRI turbulence which diffuses and damps higher radial wavenumbers. Combining a model for the damping, facilitated by the MRI, with the PI Floquet growth rates for sufficiently subsonic sloshing amplitudes does indeed suggest that there is a preference for a longer wavelength to emerge at later times. This model will be discussed in more detail in Section 5.4.
@@
Figure 12: Horizontal stresses in the equilibrium sloshing state, measured between $40<t/T<60$. $\alpha_{\rm M}$ (red line) oscillates about the averaged value of $0.0099$ (dashed red). $\alpha_{\rm R}$ (blue line) oscillates about the averaged value of $0.0019$ (dashed blue).
@@
The sloshing state also modifies the nature of the horizontal stresses, compared with the standard MRI discussed in Section 5.1. Note that when computing $\alpha_{\rm R}=\langle\rho u_{x}^{\prime}u_{y}^{\prime}\rangle_{xyz}/(q\langle\rho c^{2}\rangle_{xyz})$, the velocity fluctuations are now measured relative to the average sloshing flow. The results are shown in Fig. 12 over the time interval $40<t/T<60$, after the sloshing flow has equilibrated. The Maxwell stresses $\alpha_{\rm M}$ (red) once again dominate over the Reynolds stresses $\alpha_{\rm R}$ (blue), with time-averaged values of $0.0099$ and $0.0019$, respectively. This corresponds to an amplification of $\sim 1.4$ compared with the base state. It should also be noted that the horizontal stresses exhibit a clear oscillation over the dynamical timescale, indicating that the horizontal fluctuations are also correlated with the forcing frequency.
@@ 5.3 MHD Vertical Stresses
We will now quantify the MHD equilibrium sloshing state in detail, akin to the analysis performed in Section 4.4. In particular, we will examine the balance between the sloshing flows and the vertical stresses, which now comprise both Reynolds and Maxwell terms. The Maxwell stresses are denoted
@@
and arise from the horizontal averaging of equation (2). In this way, we extend our horizontally averaged equations for the sloshing motion by appending $\partial\mathcal{M}_{xz}/\partial z$ and $\partial\mathcal{M}_{yz}/\partial z$ to equations (27)-(28), respectively.
Figure 13: Rate of strain and stress relationship for our fiducial MHD run (akin to Fig. 9). Columns (1) and (2) respectively show the $x$ and $y$ components of the phase-folded, vertical profiles for the (a) sloshing motions, (b) Reynolds stresses, and (c) Maxwell stresses, at different orbital phases indicated by the line color. Column (3) then extracts the corresponding amplitudes of these sinusoidal profiles as a function of the temporal phase.
In Fig. 13, we compare the amplitude and phase relationship between the equilibrium sloshing motion $\mathcal{R}$ and $\mathcal{M}$. As per the analysis presented in Fig. 9, we phase fold our data modulo the dynamical period, over the interval $40<t/T<60$, to boost our sloshing signal. In panels (a1) and (a2), we are plotting the horizontally averaged sloshing motion $v_{x,{\rm w}}$ and $v_{y,{\rm w}}$, as a function of $z$, across a number of orbital phases. Similarly, we extract the phase-folded $z$-profiles for $\mathcal{R}_{xz}$, $\mathcal{R}_{yz}$, $\mathcal{M}_{xz}$, and $\mathcal{M}_{yz}$ in panels (b1), (b2), (c1), and (c2). In keeping with the hydrodynamic results, we cleanly extract oscillating sinusoidal profiles for the sloshing motion and $\cos(k_{\rm b}z)$ structures for the stresses. At each phase, we find the best fit to these profiles and thereby extract $A(t)$, $B(t)$ (panel (a3)), $R_{ij}$ (panel (b3)), and $M_{ij}$ (panel (c3)), where $\mathcal{M}_{ij}=M_{ij}(t)\cos{(k_{\rm b}z)}$.
By Fourier analyzing these curves, we isolate the amplitude and phase for the dominant orbital frequency component. We find that $|A|=0.221$ and $\arg(A)=-0.66$, whilst $|B|=0.079$ and $\arg(B)=2.06$. This shows that the sloshing amplitude is slightly larger in this case, compared with the previous hydrodynamic run. Performing the same analysis on the stresses, we find $|R_{xz}|=0.0125$ and $\arg(R_{xz})=2.98$, whilst $|R_{yz}|=0.0134$ and $\arg(R_{yz})=-1.40$. Finally, $|M_{xz}|=0.0037$ and $\arg(M_{xz})=2.04$, whilst $|M_{yz}|=0.0031$ and $\arg(M_{yz})=-1.44$. Clearly, the Reynolds stresses dominate the vertical transport, owing to the dominance of the PI in establishing strong elevator flows that resist the sloshing motions.
Once again, we can find the viscoelastic linear relationship between the velocity shear and the stresses, such that $\mathcal{R}_{ij}=\bar{\rho}\nu_{{\rm R},ij}\partial v_{i,\rm{w}}/\partial z$ and $\mathcal{M}_{ij}=\bar{\rho}\nu_{{\rm M},ij}\partial v_{i,\rm{w}}/\partial z$. In addition to equation (30), we have an analogous expression involving the Maxwell viscoelastic coefficients
@@ 5.4 Variation with Forcing Amplitude
Figure 14: Variation in the sloshing diagnostics with $\psi_{\rm f}$. Row (a): Sloshing motion $A$ and $B$. Row (b): Reynolds stresses. Row (c): Maxwell stresses. Row (d): viscoelastic constants. Column (1) visualizes the full complex information in a polar plot, whilst columns (2) and (3) plot the amplitude and phase information, respectively, as a function of $\psi_{\rm f}$. The red and black line colors denote the $x$ vs $y$ variables, whilst the different line styles+markers denote the MHD, inviscid hydrodynamic and inviscid MHD-seeded hydrodynamic runs (see inset legend). The marker colors also denote the value of $\psi_{\rm f}$. Finally, the starred markers correspond to the maximum radial sloshing amplitude attained during the transient phase after the forcing is switched on. For further explanation, see the discussion in-text.
@@
Up until now, we have focused on the fiducial forcing amplitude with $\psi_{\rm f}=0.3$. This value is sufficiently strong that the emerging sloshing equilibrium is well distinguished from the underlying MRI turbulence. However, in this section, we will consider how the sloshing response varies as a function of $\psi_{\rm f}$. To this end, we perform a suite of MHD simulations with the same setup as described in Section 5.2 but with $\psi_{\rm f}\in\{0.01,0.02,0.03,0.05,0.08,0.1,0.2,0.3\}$. In all cases, we first allow the sloshing to progress through the initial transient growth before settling into a quasi-steady oscillation. We then monitor this between times $40<t/T<60$ to extract the sloshing amplitudes, stresses, and viscoelastic coefficients, as before. The results are shown in Fig. 14. Row (a) plots the extracted sloshing amplitudes, rows (b) and (c) show the Reynolds and Maxwell stresses, respectively, and row (d) shows the viscoelastic constants. Column (1) visualizes the full complex amplitude and phase information in the form of polar plots. Meanwhile, columns (2) and (3) break this up into the amplitude and phase variation as a function of $\psi_{\rm f}$. In all cases, the colored markers denote the different values for $\psi_{\rm f}$ whilst the solid, connective colored black and red lines with circular markers pertain to the ‘$x$’ versus ‘$y$’ quantities, as denoted by the inset labels. For comparison, we also plot a sample of purely hydrodynamic runs in columns (2) and (3). Repeating the methodology described in Section 4.1 for $\psi_{\rm f}\in\{0.01,0.03,0.1,0.3\}$, we similarly extract the sloshing diagnostics between $40<t/T<60$ and plot the results as the dashed lines with triangular markers. Finally, we repeat these hydrodynamic runs but now start from the perturbed velocity and density initial condition inherited from the final snapshot of our base MRI run, with the magnetic field now switched off. These are shown as the dotted lines with square markers. It should be noted that for these hydrodynamic runs, the explicit viscosity is switched off in the simulations, in keeping with the setup described in Section 4. One might reasonably ask whether the explicit viscosity in the MHD experiments, versus the lack thereof in the hydrodynamic runs, is the main culprit behind differences in the final sloshing state. Therefore, in appendix B, we repeat the hydrodynamic runs but with an explicit viscosity that matches our MHD runs. We find essentially the same behavior as our inviscid hydrodynamic runs across the range of forcings considered. This underlines that the main conclusions of this section are insensitive to the small values of $\nu$ and it is in fact the MRI turbulence that leads to the primary differences between the MHD and hydrodynamic results.
Across all runs, panel (a2) clearly shows the unsurprising increase in the sloshing response with the forcing amplitude. However, it should be noted that this increase is sublinear in character. Indeed, fitting a power law to the MHD data yields $|A|\propto\psi_{\rm f}^{0.51}$. This owes to the fact that the viscoelastic coefficients also increase with the forcing in panel (d2). For example, the MHD runs are fit by $|\nu_{xz}|\propto\psi_{\rm f}^{0.57}$. These scalings are therefore approximately consistent with the the simplified amplitude equilibrium described by equation (17). Such fits to the viscoelastic coefficients, as a function of $\psi_{\rm f}$, therefore outline a predictive procedure wherein the sloshing response can be obtained via the laminar viscoelastic equations (15)-(16).
@@
Returning to broad trends, we see that for lower values of $\psi_{\rm f}$, the ratio of $|A|/|B|\sim 2$ whilst the phase difference in panel (a3) is also roughly $\pi/2$, in keeping with a free epicylic motion. As $\psi_{\rm f}$ increases, this ratio $|A|/|B|$ rises and the phase difference trends toward $\pi$. Comparing the various physical setups, we clearly observe differences between the corresponding line styles. Notably, the magnetized, MRI runs typically saturate with a higher sloshing amplitude compared with the hydrodynamic runs, particularly for intermediate values of $\psi_{\rm f}$. This is equivalently borne out by the magnetized runs showing lower viscoelastic amplitudes in panel (d2). This might seem surprising at first, since naively the action of the MRI might be thought to enhance the damping of the sloshing motions. However, the MRI turbulence also acts to damp the growth of the parametric instability and the associated elevator modes. Since the hydrodynamic, PI modes dominate the vertical stress, suppression of these via the MRI allows the sloshing to attain a larger amplitude. The fact that the hydrodynamic run, with the MRI inherited initial conditions, also deviates from the full MHD run, suggests that the final sloshing state is indeed sensitive to the different physics and not just the seeding of perturbations.
@@
In all panels, the vertical, dashed line and the bounded gray area, for $\psi_{\rm f}<\psi_{c}=0.03$, delimit the region for which the MRI is found to suppress the growth of the PI in our MHD runs. Namely, below this value, the runs did not show the development of elevator modes and instead the turbulence resembled the MRI base state. Meanwhile, the elevator modes and PI are still found to emerge below this value in our purely hydrodynamic runs. This MHD suppression of the PI below $\psi_{c}$ also seems to manifest in the slopes of the sloshing and viscoelastic amplitudes. The sloshing amplitude in panel (a2) rises more quickly with $\psi_{\rm f}$ when the MRI suppresses the PI, before the ‘knee’ at $\psi_{c}$, where the response flattens. The viscoelastic coefficients also seem to rise more rapidly beyond this critical forcing as the PI overcomes the MRI.
@@
A simplified toy model might be invoked to motivate this critical forcing value. In particular, the elevator modes can emerge when the PI growth rate, at the maximal value of sloshing during the initial transient phase, exceeds the damping induced by the turbulent transport of momentum across the boundary between elevator mode channels. This turbulent damping of the elevator modes can be estimated as $\sigma_{\rm d}\sim 4k_{x}v_{x,{\rm MRI}}/(2\pi)$. The factor of 4 corresponds to the momentum transport in/out of the two boundaries, either side of a vertical elevator channel. The $k_{x}=2\pi n_{x}/L_{x}$ encodes the radial width of the elevator mode, given the radial mode order $n_{x}$ fitting into the box of width $L_{x}$. Finally, the $v_{x,{\rm MRI}}$ denotes a characteristic radial velocity associated with the MRI eddies. Here, we take this to be equal to the standard deviation of the velocity fluctuation distribution, $v_{x,{\rm MRI}}=0.085$, as found in Fig. 10. This competes with the PI growth rates $\sigma_{\rm PI}$, which can be found using the Floquet theory (using $N_{\rm max}=30$), as described in Section 3.2. We solve for the maximally growing mode across a number of sloshing amplitudes $S$ and radial modes $n_{x}\leq 30$. Then, for each value of $S$, we find the maximum net growth rate $\sigma_{\rm net}=\sigma_{\rm PI}-\sigma_{\rm d}$ and the corresponding optimal modal number $n_{x}$. The results are shown in Fig. 15, where the maximal growth rate is normalized by the value of $S$.
@@
Figure 15: Model for the competition between PI growth and MRI damping of elevator modes. The maximum net growth rate $\sigma_{\rm net}$, normalized by the sloshing amplitude $S$, is plotted as a function of $S$. The color of the markers denotes the preferred radial order $n_{x}$ corresponding to this maximal growth rate.
@@
Furthermore, the color of the starred data points indicates the modal number $n_{x}$ with the largest growth rate. For $S\leq 0.1$ the MRI-induced damping suppresses the PI growth for all modes. However, above this value, our model predicts growth at $n_{x}=2$, before shorter wavelengths can overcome the damping at larger values of $S$. As one progresses to larger values of $S$, we see the preference for $n_{x}=5$, then $n_{x}=6$, in keeping with the elevator flows seen in our fiducial runs. As one progresses above $S\sim 1$ we find the optimal $n_{x}$ decreases again, in accordance with the leftward drifting peaks seen in Fig. 6.
@@
To draw a comparison with the behavior seen in our MHD simulations, we return to Fig. 14 and plot the maximum radial sloshing amplitude during the transient phase as starred markers in panel (a2). For $\psi_{\rm f}=0.03$ this yields a maximal sloshing of $|A|=0.12$, which should be susceptible to growth of the $n_{x}=2$ elevator mode according to Fig. 15. Upon examining the sloshing state for this run, we do indeed see the emergence of this very mode, marking the transition to the PI-dominated regime. Note that the maximal sloshing amplitude is always reduced substantially as the PI feeds off this shear energy and establishes stresses that equilibrate the sloshing at a lower amplitude.